OpenStudy (anonymous):

What is the solution to the system of equations? Use the linear method. 5x+2y=-4 -3x+2y=12 A. (-2,3) B. (4,0) C. (0,-2) D. (2,-7)

OpenStudy (anonymous):

@Surana

OpenStudy (anonymous):

First arrange both equations using y=mx+c

OpenStudy (anonymous):

\[5x+2y=-4-First~eqn\]First subtract both sides by 5x Then divide both sides by 2

OpenStudy (anonymous):

\[5x-5x+2y=-5x-4\]

OpenStudy (anonymous):

\[2y=-5x-4\]

OpenStudy (anonymous):

\[\frac{ 2y }{ 2 }=\frac{- 5x }{ 2 }+\frac{ -4 }{ 2 }\]

OpenStudy (anonymous):

\[y=-\frac{ 5 }{ 2 }x-2\]

OpenStudy (anonymous):

\[-3x+2y=12-Second~eqn\]

OpenStudy (anonymous):

Add 3x on both sides Then divide both sides by 2

OpenStudy (anonymous):

\[-3x+3x+2y=3x+12\]

OpenStudy (anonymous):

\[2y=3x+12\]

OpenStudy (anonymous):

\[\frac{ 2y }{ 2 }=\frac{ 3x }{ 2 }+\frac{ 12 }{ 2 }\]

OpenStudy (anonymous):

\[y=\frac{ 3 }{ 2 }x+6\]

OpenStudy (anonymous):

Now substitute the first eqn into the second eqn to find the value of x.

OpenStudy (anonymous):

\[y=-\frac{ 5 }{ 2 }x-2-First~eqn\]\[y=\frac{ 3 }{ 2 }x+6-Second~eqn\]

OpenStudy (anonymous):

\[-\frac{ 5 }{ 2 }x-2=\frac{ 3 }{ 2 }x+6\]

OpenStudy (anonymous):

First subtract 3x/2 on both sides

OpenStudy (anonymous):

\[-\frac{ 5 }{ 2 }x-\frac{ 3 }{ 2 }x-2=6\]Then add 2 on both sides

OpenStudy (anonymous):

\[-\frac{ 5 }{ 2 }x-\frac{ 3 }{ 2 }x=6+2\]

OpenStudy (anonymous):

\[\frac{ -5x-3x }{ 2 }=8\]Then multiply oth sides by 2

OpenStudy (anonymous):

both*

OpenStudy (anonymous):

\[-8x=16\]Then divide -8 on both sides

OpenStudy (anonymous):

\[x=-2\]

OpenStudy (anonymous):

Nowsubstitute x=-2 into the first equation\[y=-\frac{ 5 }{ 2 }x-2\]

OpenStudy (anonymous):

sorry I was afk

OpenStudy (anonymous):

\[y=-\frac{ 5 }{ 2 }(-2)-2\]

OpenStudy (anonymous):

\[y=5-2\]\[y=?\]solve it @mc2014

OpenStudy (anonymous):

3 So its A?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

I gave you a medal and I fanned you

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

np (no problem)